SuicideSheep32:

Raul works at a movie theatre. The function f(x) represents the amount of money Raul earns per ticket, where x is the number of tickets he sells. The function g(x) represents the number of tickets Raul sells per hour, where x is the number of hours he works. Show all work to find f(g(x)), and explain what f(g(x)) represents. f(x) = 2x^2 + 16 g(x) = sqrt 5x^3

SuicideSheep32:

@Shadow

Gdeinward:

do you know were to start?

SuicideSheep32:

honestly no

Gdeinward:

how confortable with algebra are you?

Gdeinward:

I just want to make sure you are aware how to solve functions and such

SuicideSheep32:

i understand most things but square roots mess me up

Gdeinward:

okay cool

Gdeinward:

essentially f(g(x)) means you need to input g(x) into the x terms of th f(x) equation

SuicideSheep32:

ok so f(g(x)) = 2(sqrt 5x^3)^2+16 ?

Gdeinward:

that is correct

SuicideSheep32:

ok so what do i do to solve

SuicideSheep32:

more like where do i start?

Gdeinward:

you need to distribute the parentesis

Gdeinward:

so the new equation would be 2(squrt5^3)(squrt5^3)+ 16

Gdeinward:

you know what im gonna invite someone over to help. Im in the middle of history so my mind isnt right for math.

Gdeinward:

@mhchen

SuicideSheep32:

sorry I was working on another question while i waited and got carried away

mhchen:

is \[g(x) = \sqrt{5x^{3}}\] ?

SuicideSheep32:

Yes

mhchen:

so \[f(g(x)) = 2(\sqrt{5x^{3}})^{2}+16\] right?

SuicideSheep32:

yes that is what I got so far

mhchen:

Then \[f(g(x)) = 2(5x^{3})+16\] right?

SuicideSheep32:

ok i get that

mhchen:

Then \[f(g(x)) = 10x^{3}+16\] right?

SuicideSheep32:

yes ok

mhchen:

Yeah...that's the most you can do. So "f(x) represents the amount of money Raul earns per ticket, where x is the number of tickets he sells" and that means Raul's money = x ticket sold. "g(x) represents the number of tickets Raul sells per hour, where x is the number of hours he works" and that means tickets sold = x hours. right?

SuicideSheep32:

thank you so much you really did help me.